OpenStudy (twizttiez):

A machine coats metal spheres with a plastic coating. The metal spheres have a radius of 9 mm. The spheres formed by the metal and plastic coating have a radius of 12 mm. How much plastic coating is on each sphere? Use 3.14 to approximate pi and express your final answer in hundredths. mm3

OpenStudy (twizttiez):

@死は最高です

OpenStudy (cookie_2046):

Find the volumes of the two spheres and subtract. V = 4/3 * pi * r^3 V = 4/3 * 3.14 * 9^3 Simplify exponent: V = 4/3 * 3.14 * 729 Multiply: V = 3052.08 V = 4/3 * pi * r^3 V = 4/3 * 3.14 * 12^3 Simplify exponent: V = 4/3 * 3.14 * 1728 Multiply: V = 7234.56 Subtract the two volumes: 7234.56 - 3052.08 = 4182.48

OpenStudy (twizttiez):

Thank you

OpenStudy (cookie_2046):

You're Welcome :)